3. a) If the solution of the recurrence relation αun−1 +βun−2 = f(n),(n ≥ 2) is

un = 1−2n+3.2

n

, then determine the values of α,β and f(n).


1
Expert's answer
2021-04-15T07:28:01-0400

Given recurrence relation is-

"\u03b1u_{n\u22121} +\u03b2u_{n\u22122} = f(n)"


Also, "u_n=1-2n+3.2^n"


For solving such equation the value of coefficient is 1 so "\alpha=1,\beta=1"

The solution of the above equation can be written by characterstics root method as-

"u_n=c_1+c_2n+3.2^n"


The value of f(n) must be the value of the nin homogeneous part of the solution i.e. "f(n)=3.2^n"